OpenStudy (anonymous):

Zeus is 20 years older than Athena. In two years he will be twice her age. How old are they now?

OpenStudy (anonymous):

zeus is 38, athena is 18. can i have your phone number?

OpenStudy (anonymous):

how did you solve it? @ryandaman2017

OpenStudy (kendricklamar2014):

38 - 18 = 20 38 + 2 = 40, 18 + 2 = 20 20 * 2 = 40

OpenStudy (anonymous):

Text me an ill show you how i did? @ninamae05